For a mean of 75 and a standard deviation of 3, find P(X ≥ 77 )
1
Expert's answer
2020-06-01T17:59:04-0400

When the distribution is not mentioned, it is conventionally assumed to be Normal.

Here the distribution of X has not been mentioned, so we assume

X ~ N("\mu" = 75, "\sigma"2 = 32)


Then by the property of Normal distribution,


Z = "\frac{x-\mu}{\sigma}" ~ N(0, 1) i.e. Z is a standard normal variable


Therefore, P(X "\geq" 77)


= P("\frac{x-75}{3}\geq\frac{77-75}{3}")


= P(Z "\geq" 0.67)


=1 - P(Z < 0.67)


= 1 - "\Phi"(0.67)


= 1 - 0.7486 = 0.2514 [using standard normal distribution table]


Answer: The value of P(X ≥ 77) is 0.2514.